Probability

A recent survey reported in BusinessWeek dealt with the salaries of CEOs at large corporations and whether company shareholders made money or lost money.

CEO Paid
More Than
$1 Million CEO Paid
Less Than
$1 Million Total
Shareholders made money 1 13 14
Shareholders lost money 6 3 9

Total 7 16 23


If a company is randomly selected from the list of 23 studied, calculate the probabilities for the following :

(a) The CEO made more than $1 million. (Round your answer to 2 decimal places.)

Probability

(b)

The CEO made more than $1 million or the shareholders lost money. (Round your answer to 2 decimal places.)

Probability

(c)

The CEO made more than $1 million given the shareholders lost money. (Round your answer to 4 decimal places.)

Probability

(d)

Select two CEOs and find that they both made more than $1 million. (Round your answer to 4 decimal places.)

Probability
Answer #1
(a) The CEO made more than $1 million. (Round your answer to 2 decimal places.)

  Probability = \frac{7}{23} = 0.3043

--------------------------------------------------------------------------------------------

b)

The CEO made more than $1 million or the shareholders lost money. (Round your answer to 2 decimal places.)

  Probability = 7+9-6/23 = \frac{10}{23} = 0.4348

---------------------------------------------------------------------------------------

Probability   
(c)

The CEO made more than $1 million given the shareholders lost money. (Round your answer to 4 decimal places.)

  Probability

= \frac{6}{23} divided by \frac{9}{23} = \frac{6}{9} =\frac{2}{3} = 0.6667

--------------------------------------------------------

  Probability   
(d)

Select two CEOs and find that they both made more than $1 million. (Round your answer to 4 decimal places.)

We have to find probability that 2 ceos made more than 1 million.

p =7/23 and q=16/23.

  Probability = 7C2 (7/23)^2 (16/23)^5   = 0.3169
Answer #2

The CEO made more than $1 million = 7/23

The CEO made more than $1 million or the shareholders lost money. =6/7

The CEO made more than $1 million given the shareholders lost money. = 6/7 (9/23) =33.54%

Select two CEOs and find that they both made more than $1 million. = (7/23)2 =.0926 0r 9.26%.